Required information [The following information applies to the questions displayed below.] Corrigan Enterprises is studying the acquisition of two electrical component insertion systems for producing its sole product, the universal gismo. Data relevant to the systems follow. Model no. 6754 Variable costs, $17.00 per unit Annual fixed costs, $986,100 Model no. 4399: Variable costs, $11.80 per unit Annual fixed costs, $1,114,100 Corrigan's selling price is $61 per unit for the universal gismo, which is subject to a 15 percent sales commission. (In the following requirements, ignore income taxes.) Required: 1. How many units must the company sell to break even if Model 6754 is selected? (Do not round intermediate calculations and round your final answer up to nearest whole number.) Break-even point units
2-a. Calculate the net income of the two systems if sales and production are expected to average 43,000 units per year. Net Income Model No. 6754 Model No. 4399 2-b. Which of the two systems would be more profitable? Model No. 6754 Model No. 4399
3. Assume Model 4399 requires the purchase of additional equipment that is not reflected in the preceding figures. The equipment will cost $440,000 and will be depreciated over a five-year life by the straight-line method. How many units must Corrigan sell to earn $969,000 of income if Model 4399 is selected? As in requirement (2), sales and production are expected to average 43,000 units per year. (Do not round intermediate calculations and round your final answer up to nearest whole number.) units Required sales
4. Ignoring the information presented in part (3), at what volume level will the annual total cost of each system be equal? (Do not round intermediate calculations and round your final answer up to nearest whole number.) Volume level units
Answer #1

1.Break even Point –Model 6754

ANs:

28,296 Units

Break Even Point=Fixed cost expenses/Contribution Margin

=$986,100/$34.85=282,95.55=28,296 Units (Rounded off)

Contribution Margin=Sales-Variable Cost-other variable cost(commission)

=$61-$17-$9.15(61*15%)=$34.85

2.a Net income (model no:6754)

Sales revenue (43000*61)                                                            $ 2623,000

Less: variable cost (43000*17)                    ($731,000)

       :other variable expenses                       ($393,450)

      (sales commission:2623,000*15%)

     Total Variable cost                                                                      ($1124,450)

Contribution Margin                                                                       $1498,550

Less: fixed Cost                                                                                 (986,100)

Net Income                                                                                        512,450

.a Net income (model no:4399)

Sales revenue (43000*61)                                                            $ 2623,000

Less: variable cost (43000*11.80)              ($507,400)

       :other variable expenses                       ($393,450)

      (sales commission:2623,000*15%)

     Total Variable cost                                                                      ($900,850)

Contribution Margin                                                                       $1722150

Less: fixed Cost                                                                                 ($1114,100)

Net Income                                                                                        $608,050

2.b.

Ans; Model No 4399

3.

Compute target sales to earn desired profit

Target sales=Fixed cost+Target Profit/Contribution

=1202100+969,000/40.05=54209.73=54,210

The sales units are 54,210 units

Workings:

Computation of fixed cost for model no 4399    1114,100

Add: Depreciation($440,000/5) 88,000

Total Fixed Cost                                                                1202100                               

Contribution margin per unit= Sales-Variable Cost-other variable cost(commission)

$61-$11.80-$9.15(61*15%)=$40.05

4. Indifferent point

Total annual cost of Model No 6754 = Total annual cost of model no 4399

17X+9.15X+986100 = 11.80X+9.15X+1114,100

26.15X-20.95X = 1114,100-986100

5.2X = 128,000

X(volume) = 128,000/5.2 = 24,615 Units